OpenStudy (anonymous):

Match the factorial to its value. 1. 7! 40,320 2. 5! 720 3. 8! 120 4. 6! 5,040

OpenStudy (anonymous):

7! = 1 * 2 * 3 * 4 * 5 * 6 * 7 = ?! 6! = 1 * 2 * 3 * 4 * 5 * 6 = !? and so on

OpenStudy (anonymous):

@chrisneedshelp avavavaveee

OpenStudy (anonymous):

1!<2!<3!<4!<5!.... don't need to calculate the values just sort them from lowest to highest and sort the factorial base from lowest to highest and then match the positions to each other